OpenStudy (anonymous):

What is the point of symmetry for the circle (x + 2)2 + (y – 4)2 = 25?

OpenStudy (anonymous):

find intercepts

OpenStudy (anonymous):

x=0 => y=3, -3

OpenStudy (anonymous):

no y =3,-3 sorry

OpenStudy (anonymous):

Equation for a circle, with center at (a,b): \[(x-a)^2 + (y-b)^2 = radius^2\] Can you find the center? :-)

OpenStudy (anonymous):

The point of symmetry is the center. \[(x - h)^{2} + (y - k)^{2} = r^{2}\]\[(x + 2)^{2} + (y - 4)^{2} = r^{2}\] Where the center is (-h, -k)

OpenStudy (anonymous):